A car going 80 mph passes a car setting still after the car passes the car travels 100 mph to catch the other car how long will it take the car travel 100 mph to catch up to the car travel 80 mph
1
Expert's answer
2012-06-01T07:17:38-0400
Let the time before the second car will start ts. Time to catch up the
first
car tc.
80 * ts = 80 * tc + 100 * tc
tc = (80 * ts) / (80 + 100)
tc =
0,4444 * ts